OpenStudy (anonymous):

what is 11/12 - 5/12? 3/4 1/2 2/3 6/0

jhonyy9 (jhonyy9):

6/12 =1/2

OpenStudy (anonymous):

6/12 = 1/2

OpenStudy (anonymous):

ahhh second too late

OpenStudy (anonymous):

ok thnx :))))))

jhonyy9 (jhonyy9):

yw